# Nonasymptotic Analysis of Stochastic Gradient Descent with the Richardson–Romberg Extrapolation

### v2

@@ A.3 Proof of Section 4
$\|\mathsf{E}_{\nu}[\bar{\theta}_{n}^{(RR)}-\theta^{\star}]\|\leq(\frac{L}{\mu}\mathsf{C}_{1}+\frac{L\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{5/2}})\gamma^{3/2}+\mathcal{R}_{3}(\theta_{0}-\theta^{\star},\gamma,n),$ (63)
@@ Appendix B Proof of Theorem 1
Assume A 1, A 2, A 3($6$), and C 1($6$). Then for any $\gamma\in(0,1/(\operatorname{L}\mathsf{C}_{\operatorname{step},6})]$, $n\in\mathbb{N}$, and initial distribution $\nu$ on $\mathbb{R}^{d}$, the sequence of Polyak-Ruppert estimates (14) satisfies
@@
[…]
@@ Appendix C Proof of Theorem 2
$\begin{split}\mathsf{C}_{\operatorname{RR},1}&=\frac{c_{0}\mathsf{D}_{\operatorname{last},4}\operatorname{L}\tau_{4}^{2}}{\mu^{3/2}}+\frac{c_{0}\operatorname{L}\mathsf{D}_{\operatorname{last},2}^{1/2}\tau_{2}}{\mu^{1/2}}\,,\quad\mathsf{C}_{\operatorname{RR},2}=\frac{c_{0}\mathsf{D}_{\operatorname{last},2}^{1/2}\tau_{2}}{\mu^{1/2}}\\ \mathsf{C}_{\operatorname{RR},3}&=c_{0}\left(\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{\mu^{3/2}}+\mathsf{C}_{1}\right)\,,\quad\mathsf{C}_{\operatorname{RR},4}=\frac{c_{0}\mathsf{D}_{\operatorname{last},4}\operatorname{L}\tau_{4}^{2}}{\mu}\,,\end{split}$ (73)
$\mathsf{C}_{1}$ is defined in (51), and the remainder term $\mathcal{R}_{4}(n,\gamma,\|\theta_{0}-\theta^{\star}\|)$ is given by
@@
[…]
@@
$\mathsf{E}_{\pi_{\gamma}}[(\psi(\theta_{0})-\pi_{\gamma}(\psi))^{T}(\psi(\theta_{k})-\pi_{\gamma}(\psi))]$ (79)
$\qquad=\mathsf{E}_{\pi_{\gamma}}[(\psi(\theta_{0})-\pi_{\gamma}(\psi))^{T}({\rm Q}_{\gamma}^{k}\psi(\theta_{0})-\pi_{\gamma}(\psi))]$ (80)
@@
[…]
@@
[…]
@@
Using A 1 and since $41-\gamma^{3}\operatorname{L}^{3}/4-3\gamma\operatorname{L}/2-\gamma^{2}\operatorname{L}^{2}\geq 1-11\gamma\operatorname{L}/$ 4, we get
@@
Since $21-11\gamma\operatorname{L}/4\geq 1/$ 2 for $\gamma\leq 2/(11\operatorname{L})$, we complete the proof.
@@
[…]
@@
[…]
@@
[…]
@@ Appendix D Proof of Theorem 3
$\begin{split}\mathsf{C}_{\operatorname{RR},5}&=\frac{c_{0}\mathsf{D}_{\operatorname{last},p}^{1/2}\tau_{p}}{\mu^{1/2}}\,,\quad\mathsf{C}_{\operatorname{RR},6}=\frac{c_{0}\operatorname{L}\mathsf{D}_{\operatorname{last},p}^{1/2}p\tau_{p}}{\mu^{1/2}}+\frac{c_{0}\operatorname{L}\mathsf{D}_{\operatorname{last},2p}p\tau^{2}_{2p}}{\mu^{3/2}}\,,\\ \mathsf{C}_{\operatorname{RR},7}&=c_{0}\left(\mathsf{C}_{1}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},3p}^{3/2}\tau_{3p}^{3}}{\mu^{3/2}}\right)\,,\quad\mathsf{C}_{\operatorname{RR},8}=\frac{c_{0}\operatorname{L}\mathsf{D}_{\operatorname{last},2p}\tau_{2p}}{\mu^{2}}\,,\end{split}$ (101)
$\mathsf{C}_{1}$ is defined in (51), and the remainder term $\mathcal{R}_{5}(n,\gamma,\|\theta_{0}-\theta^{\star}\|)$ is given by
@@ D.1 Proof of Section 4
$\|\mathrm{Q}^{mk}\psi(\theta)-\pi_{\gamma}(\psi)\|\leq\frac{1}{2}\operatorname{L}_{3}\mathbf{W}_{{c}}(\delta_{\theta}{\rm Q}_{\gamma}^{km},\pi_{\gamma})\leq 2\operatorname{L}_{3}(1/2)^{k}\mathbf{W}_{{c}}(\delta_{\theta},\pi_{\gamma})\,,$
@@
In order to prove Theorem 6, we need a generalization of Section 4 for arbitrary initial distribution $\nu$. We provide this result below.
[…]
Under assumptions of Section 4 for any $\gamma\in(0,1/(L\mathsf{C}_{\operatorname{step},6})],n\in\mathbb{N}$ and initial distribution $\nu$, it holds that
$\mathsf{E}^{1/p}_{\nu}\bigl{[}\|\sum_{k=n+1}^{2n}\{\psi(\theta_{k})-\pi_{\gamma}(\psi)\}\|^{p}\bigr{]}\leq\mathsf{E}^{1/p}_{\pi_{\gamma}}\bigl{[}\|\sum_{k=n+1}^{2n}\{\psi(\tilde{\theta}_{k})-\pi_{\gamma}(\psi)\}\|^{p}\bigr{]}\\ +\frac{\operatorname{L}_{3}(1-\gamma\mu)^{(n+1)/2}}{\gamma\mu}\left(\mathsf{E}^{1/p}_{\nu}[\|\theta_{0}-\theta^{\star}\|^{2p}]+\frac{\mathsf{D}_{\operatorname{last},2p}\gamma\tau_{2p}^{2}}{\mu}\right)\,.$
@@
Using Hölder’s and Minkowski’s inequalities together with C 1($2p$) and Appendix C, we obtain
$(\mathsf{E}^{{\rm K}_{\gamma}}_{\nu,\pi_{\gamma}}[c^{p}(\theta_{k},\tilde{\theta}_{k})])^{1/p}$
@@
$\biggl{(}\mathsf{E}^{{\rm K}_{\gamma}}_{\nu,\pi_{\gamma}}[\|\sum_{k=n+1}^{2n}\{\psi(\theta_{k})-\psi(\theta_{k}^{\prime})\}\|^{p}]\biggr{)}^{1/p}\lesssim\frac{\operatorname{L}_{3}(1-\gamma\mu)^{(n+1)/2}}{\gamma\mu}\biggl{(}\mathsf{E}^{1/p}_{\nu}[\|\theta_{0}-\theta^{\star}\|^{2p}]+\frac{\mathsf{D}_{\operatorname{last},2p}\gamma\tau_{2p}^{2}}{\mu}\biggr{)}\,,$

### v3

@@ A.3 Proof of Section 4
$\|\mathsf{E}_{\nu}[\bar{\theta}_{n}^{(RR)}-\theta^{\star}]\|\leq(\frac{L}{\mu}\mathsf{C}_{b,2}+\frac{L\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{5/2}})\gamma^{3/2}+\mathcal{R}_{3}(\theta_{0}-\theta^{\star},\gamma,n),$ (64)
@@ Appendix B Proof of Theorem 3
Assume A 1, A 2, A 3($4$), and C 1($4$). Then for any $\gamma\in(0,1/(\operatorname{L}\mathsf{C}_{\operatorname{step},4})]$, $n\in\mathbb{N}$, and initial distribution $\nu$ on $\mathbb{R}^{d}$, the sequence of Polyak-Ruppert estimates (14) satisfies
@@
$\nabla f(\theta_{k})=\operatorname{H}^{\star}(\theta_{k}-\theta^{\star})+\left(\int_{0}^{1}t\nabla^{3}f(t\theta^{\star}+(1-t)\theta_{k})\,dt\right)(\theta_{k}-\theta^{\star})^{\otimes 2}\,$,
@@ Appendix C Proof of Theorem 6
[…]
and the remainder term $\mathcal{R}_{4}(n,\gamma,\|\theta_{0}-\theta^{\star}\|)$ is given by
@@
$G(\theta)=\frac{1}{2}\left(\int_{0}^{1}t^{2}\nabla^{4}f(t\theta^{\star}+(1-t)\theta)\,dt\right)(\theta-\theta^{\star})^{\otimes 3}\,.$
@@
$\mathbf{W}_{{c}}(\delta_{\theta_{0}},\pi_{\gamma})=\int c(\theta_{0},\vartheta)\mathrm{d}\pi_{\gamma}(\vartheta)\,.$
@@
$\int\int c^{2}(\theta_{0},\vartheta)d\pi_{\gamma}(\vartheta)d\pi_{\gamma}(\theta_{0})$ (82)
$\leq\int\int\|\theta_{0}-\vartheta\|^{2}\left(\|\theta_{0}-\theta^{\star}\|+\|\vartheta-\theta^{\star}\|+\frac{2^{3/2}\gamma^{1/2}\tau_{2}}{\mu^{1/2}}\right)^{2}d\pi_{\gamma}(\vartheta)d\pi_{\gamma}(\theta_{0})$ (83)
@@
Using A 1 and since $1-\gamma^{3}\operatorname{L}^{3}/4-3\gamma\operatorname{L}/2-\gamma^{2}\operatorname{L}^{2}\geq 1-11\gamma\operatorname{L}/4$, we get
@@
Since $1-11\gamma\operatorname{L}/4\geq 1/2$ for $\gamma\leq 2/(11\operatorname{L})$, we complete the proof.
@@
$\eta(\theta)=\psi(\theta)+\frac{1}{2}\left(\int_{0}^{1}t^{2}\nabla^{4}f(t\theta^{\star}+(1-t)\theta)\,dt\right)(\theta-\theta^{\star})^{\otimes 3}\,$, (92)
@@
$(1/2)\|\left(\int_{0}^{1}t^{2}\nabla^{4}f(t\theta^{\star}+(1-t)\theta)\,dt\right)(\theta-\theta^{\star})^{\otimes 3}\|\leq\operatorname{L}_{4}\|\theta-\theta^{\star}\|^{3}\,.$ (93)
@@
$\|T\|=\underset{x\neq 0,y\neq 0,z\neq 0}{\sup}\frac{\underset{i,j,k}{\sum}{T_{ijk}x_{i}y_{j}z_{k}}}{\|x\|\|y\|\|z\|}\geq\underset{x\neq 0,y\neq 0}{\sup}\underset{z\neq 0}{\sup}\frac{\underset{k}{\sum}z_{k}\underset{i,j}{\sum}T_{ijk}x_{i}y_{j}}{\|z\|\|y\|\|x\|}=\underset{x\neq 0,y\neq 0}{\sup}\frac{\|t(x,y)\|}{\|y\|\|x\|},$ (97)
@@ Appendix D Proof of Theorem 9
[…]
and the remainder term $\mathcal{R}_{5}(n,\gamma,\|\theta_{0}-\theta^{\star}\|)$ is given by
@@ D.1 Proof of Section 4
$\|\mathrm{Q}_{\gamma}^{mk}\psi(\theta)-\pi_{\gamma}(\psi)\|$ $=\sup_{u\in\mathbb{R}^{d}:\|u\|=1}|\mathrm{Q}_{\gamma}^{mk}(u^{\top}\psi(\theta))-\pi_{\gamma}(u^{\top}\psi)|$
$\leq\frac{1}{2}\operatorname{L}_{3}\mathbf{W}_{{c}}(\delta_{\theta}{\rm Q}_{\gamma}^{km},\pi_{\gamma})$
$\leq 2\operatorname{L}_{3}(1/2)^{k}\mathbf{W}_{{c}}(\delta_{\theta},\pi_{\gamma})\,$,
@@
In order to prove Theorem 23, we need a generalization of Section 4 for arbitrary initial distribution $\nu$. The following result holds:
Lemma 24.
Let $\{\tilde{\theta}^{(\gamma)}_{k}\}_{k\in\mathbb{N}}$ and $\{\theta^{(\gamma)}_{k}\}_{k\in\mathbb{N}}$ be defined by the synchronous coupling construction (47), where $\tilde{\theta}^{(\gamma)}_{0}\sim\pi_{\gamma}$ and $\theta^{(\gamma)}_{0}\sim\nu$. Then, under assumptions of Section 4, for any $\gamma\in(0,1/(L\mathsf{C}_{\operatorname{step},6})],n\in\mathbb{N}$ and initial distribution $\nu$, it holds that
$\mathsf{E}^{1/p}_{\nu}\bigl{[}\|\sum_{k=n+1}^{2n}\{\psi(\theta_{k}^{(\gamma)})-\pi_{\gamma}(\psi)\}\|^{p}\bigr{]}\leq\mathsf{E}^{1/p}_{\pi_{\gamma}}\bigl{[}\|\sum_{k=n+1}^{2n}\{\psi(\tilde{\theta}^{(\gamma)}_{k})-\pi_{\gamma}(\psi)\}\|^{p}\bigr{]}\\ +\frac{c_{0}\operatorname{L}_{3}(1-\gamma\mu)^{(n+1)/2}}{\gamma\mu}\left(\mathsf{E}^{1/p}_{\nu}[\|\theta^{(\gamma)}_{0}-\theta^{\star}\|^{2p}]+\frac{\mathsf{D}_{\operatorname{last},2p}\gamma\tau_{2p}^{2}}{\mu}\right)\,$,
where $c_{0}$ is an absolute constant.
@@
Using Hölder’s and Minkowski’s inequalities together with C 1($2p$) and Appendix C, we obtain that
$\bigl{(}\mathsf{E}^{{\rm K}_{\gamma}}_{\nu,\pi_{\gamma}}[c^{p}(\theta_{k},\tilde{\theta}_{k})]\bigr{)}^{1/p}$
@@
$\biggl{(}\mathsf{E}^{{\rm K}_{\gamma}}_{\nu,\pi_{\gamma}}[\|\sum_{k=n+1}^{2n}\{\psi(\theta_{k})-\psi(\tilde{\theta}_{k})\}\|^{p}]\biggr{)}^{1/p}\lesssim\frac{\operatorname{L}_{3}(1-\gamma\mu)^{(n+1)/2}}{\gamma\mu}\biggl{(}\mathsf{E}^{1/p}_{\nu}[\|\theta_{0}-\theta^{\star}\|^{2p}]+\frac{\mathsf{D}_{\operatorname{last},2p}\gamma\tau_{2p}^{2}}{\mu}\biggr{)}\,$,